OpenStudy (anonymous):

help!!!!!!!!!!!!!! -3x-(4x-6)3²-x=2x+4 (x-2) + 3²

OpenStudy (anonymous):

I got x=1 in the end.. Am I correct?

OpenStudy (anonymous):

or did I make a mistake?

OpenStudy (anonymous):

First you do the exponents and brackets first: -3x-(4x-6)9-x=2x+4x-8+9 Then you do the remaining braket: -3x-36x+54-x=2x+4x-8+9 Now you arrange them according to unknowns and numbers: -3x-36x-x-2x-4x=-54-8+9 And then you just solve: -46x=53 and you need to make the unknown by itself so: x= -53/46 or x= 1 7/46